Using Table 2.1 Convolve the following X(t) = 15e-3tu(t) h(t) = u(t) + 3e-9tu(t)
Answer #1

ct) hCt)= g(t) Ycs)xC) HCs) -3t Ciiven, ct)= 15e uct) hct) ucG+3e ucD Ond XCS) SH3 5 3 SP2 ( ST9 IS 45 s(st) 15Csta)t 45s sssCUbere (S135 A = |35 A (60s+135) &(s+9) s=-3 (GOC-15 C-3) -3+9 Go t 125) C = sCst?) L405 -x €9+3) 5/1 YCs) +9) CS+3 trans fom